OpenStudy (anonymous):

Solve the following system of equations. 2x + 3y - z = 13x + y + 2z = 12x + 2y - 3z = -5 Answer options: (3, 1, 2) (-3, 1, 2) (3, -1, 2) (3, 1, -2)

OpenStudy (shubhamsrg):

why dont you go for option checking? ;)

OpenStudy (anonymous):

I'm checking my answer momentarily

OpenStudy (anonymous):

Broke down it to 3 equations and do elimination.. :)

OpenStudy (anonymous):

2x + 3y - z = 1 3x + y + 2z = 1 2x + 2y - 3z = -5

OpenStudy (anonymous):

It's either B or C from what I can tell